OpenStudy (wintersuntime):

Solve this equation in substituion method 3x+(-3x+9)

OpenStudy (anonymous):

Hey winter

OpenStudy (wintersuntime):

hi

OpenStudy (mathmale):

Let's focus on the math, OK? What does the substitution method have to do with this problem? Was there another equation that you could use as a substitution?

OpenStudy (wintersuntime):

Equal values method

OpenStudy (mathmale):

Could you explain that? What is the goal of this problem?

OpenStudy (wintersuntime):

Substitution or equal values ?

OpenStudy (mathmale):

I'm asking you. Hard to "solve" a problem until we know where we're going. What is the goal of this problem you've posted?

OpenStudy (wintersuntime):

To solve it and finding the answer

OpenStudy (wintersuntime):

I think the first thing is to distribute ?

OpenStudy (mathmale):

Try removing the parentheses. Type out what you get by doing that. No, there's nothing to distribute here (other than the + sign).

OpenStudy (wintersuntime):

3x+-3x+9

OpenStudy (mathmale):

Must combine +-. Which sign do you mean?

OpenStudy (wintersuntime):

But don't you need the parentheses for substitution

OpenStudy (wintersuntime):

both

OpenStudy (anonymous):

So u mean plus a negative

OpenStudy (wintersuntime):

its -3 not subtracting

OpenStudy (mathmale):

Really, substitution has nothing to do with this problem, unless you've copied it incorrectly.

OpenStudy (mathmale):

So your 3x - 3x + 9 = ?

OpenStudy (wintersuntime):

In the question it says solve this system algebraically using the substiution method

OpenStudy (anonymous):

Screen shot and send plz

OpenStudy (mathmale):

You have listed only ONE equation. Note that "a system of linear equations" involves TWO linear equations. Where's the other one?

OpenStudy (wintersuntime):

y=3x and y= -3x+9

OpenStudy (mathmale):

Here you presented only 1 equation, whereas you started with 2 equations. That's where the confusion started. How would you use substitution to solve this system of linear equations?

OpenStudy (wintersuntime):

plug in 3x into the second equation where y is

OpenStudy (wintersuntime):

So it would be -3x+9(3x) ??

OpenStudy (mathmale):

All right, then. We then get 3x=-3x+9. No parentheses, notice? How would you solve this? First step?

OpenStudy (wintersuntime):

oh

OpenStudy (wintersuntime):

subtract -3x on both sides right ?

OpenStudy (mathmale):

Why subtract? What's your goal here?

OpenStudy (wintersuntime):

Its the inverse operation of adding

OpenStudy (mathmale):

Your goal is the inverse operation of adding? ???

OpenStudy (wintersuntime):

no to solve the equation

OpenStudy (mathmale):

Why would you add or subtract something to both sides of this equation? What are y ou trying to do? 3x=-3x+9 What next, and why?

OpenStudy (wintersuntime):

Add both sides ??

OpenStudy (wintersuntime):

your confusing me

OpenStudy (anonymous):

Yes

rebeccaxhawaii (rebeccaxhawaii):

@mathmale is here you are in good hands

OpenStudy (wintersuntime):

oh I got it I already solved it but I don't know if my answer is right ?

OpenStudy (wintersuntime):

x=3/2

OpenStudy (mathmale):

mathmale Medals 0 Look: You have 3x=-3x+9. Your goal is to solve this equation for x. You have x on both sides of this equation. Not good. How do you move the -3x on the right side over to the left side? Show your work. 3x=-3x+9 What next?

OpenStudy (wintersuntime):

y=9/2

OpenStudy (wintersuntime):

Is that right ??

OpenStudy (mathmale):

Explain. My question for you, Winter, was how could you get rid of that -3x on the right side and have all x terms on the left side?

OpenStudy (anonymous):

Screen shot and send plz

OpenStudy (mathmale):

If I gave you -3x, what operation, involving what quantity, would give you a zero?

OpenStudy (anonymous):

Screen shot and send plz

OpenStudy (mathmale):

In different words: You want to group ALL x terms on the left side. That's what we're trying to do here.

OpenStudy (wintersuntime):

so my answers are wrong?

OpenStudy (mathmale):

Please follow my directions.

OpenStudy (wintersuntime):

okay

OpenStudy (mathmale):

You have 3x on the left side of your equation and you have -3x on the right side. Your task is to consolidate the x terms on the LEFT side. I asked you before: Suppose I gave you -3x...What math operation would result in that -3x becoming zero?

OpenStudy (wintersuntime):

subtracting ?

OpenStudy (mathmale):

+ and - are opposite operations. If I give you -3x, you could turn that into a zero by adding 3x. Do that now, please. -3x+3x=?

OpenStudy (mathmale):

3x -3x ----- ?

OpenStudy (wintersuntime):

-3x+3x=0

OpenStudy (wintersuntime):

like terms

OpenStudy (mathmale):

Yes. That's how you get rid of the -3x on the right side. Now the entire equation is 3x=-3x+9. Solve for x. Move that -3x on the right side to the left side of this equation. You now know how to do this.

OpenStudy (wintersuntime):

3x = -3x + 9 3x = 9 + -3x 3x + 3x = 9 + -3x + 3x Combine like terms: 3x + 3x = 6x 6x = 9 + -3x + 3x Combine like terms: -3x + 3x = 0 6x = 9 + 0 6x = 9 Divide each side by '6'. x = 1.5

OpenStudy (mathmale):

Very nice. Any questions?

OpenStudy (wintersuntime):

no thank you so much

OpenStudy (wintersuntime):

I don't plug in for y??